OpenStudy (anonymous):

how do I rewrite the expression sin(pie/2 + x) in terms of sin x and cos x? thanks

OpenStudy (anonymous):

Well at x = 0 the expression will have the value \(sin(\pi/2) = 1\). The value of the expression will always ahead of the typical sin function. At \(x = \pi\) the value of the expression will be -1. We can see that these values directly correspond to the values of the cosine function at 0, and pi respectively. \(sin(\pi/2 + x) = cos(x)\)